# A Survey on Federated Fine-Tuning of Large Language Models

### v2

@@ 5.1.3 Medical Domain Instruction Fine-Tuning Datasets
For English-language datasets, ChatDoctor-200K (Li et al., 2023i) and ChatDoctor-HealthCareMagic-100k (Li et al., 2023i) contain medical dialogues derived from professional consultation platforms, facilitating multi-turn reasoning and symptom analysis. The Medical Meadow (Han et al., 2023) suite offers curated datasets from sources such as CORD-19999https://huggingface.co/datasets/medalpaca/medical__\__meadow__\__cord19 and MedQA [^10] https://huggingface.co/datasets/medalpaca/medical__\__meadow__\__medqa, targeting tasks like biomedical question answering, literature summarization, and clinical fact verification. Additionally, HealthCareMagic-100k-en [^11] https://huggingface.co/datasets/wangrongsheng/HealthCareMagic-100k-en supports English patient–doctor interactions across various medical specialties.
For Chinese medical instruction tuning, a rich set of datasets has been developed to address the unique linguistic and clinical characteristics of Chinese healthcare scenarios. Notable examples include ChatMed-Consult-Dataset (Zhu et al., 2023b) and CMtMedQA (Yang et al., 2024a), which focus on consultation-style QA pairs. DISC-Med-SFT (Bao et al., 2023), Huatuo-26M (Li et al., 2023d), Huatuo26M-Lite [^12] https://huggingface.co/datasets/FreedomIntelligence/Huatuo26M-Lite, and HuatuoGPT-sft-data-v1131313https://huggingface.co/datasets/FreedomIntelligence/HuatuoGPT-sft-data-v1 offer large-scale instruction–response pairs in clinical medicine, public health, and disease treatment. Traditional Chinese Medicine (TCM) is also covered through datasets like ShenNong-TCM-Dataset (Zhu & Wang, 2023), enabling LLMs to support specific diagnostic and treatment tasks.
@@ 5.1.4 Code Domain Instruction Fine-Tuning Datasets
Representative datasets include CodeAlpaca (Chaudhary, 2023) and Code Instructions 120k Alpaca [^14] https://huggingface.co/datasets/iamtarun/code__\__instructions__\__120k__\__alpaca, which extend the Alpaca instruction format to software engineering tasks such as debugging, function generation, and refactoring. CodeContests (Li et al., 2022c) focuses on competitive programming tasks and provides instruction–response pairs related to algorithmic problem solving. CommitPackFT (Muennighoff et al., 2023) offers commit-message generation tasks based on source code diffs, reflecting real-world software maintenance scenarios. ToolBench (Qin et al., 2023a) is designed to help models learn tool-augmented code generation through instruction-following examples.
Several large-scale pretraining and fine-tuning datasets are also widely used for code instruction alignment. CodeParrot-Clean [^15] https://huggingface.co/datasets/codeparrot/codeparrot-clean and The Stack v2 Dedup (Kocetkov et al., 2022) provide diverse and deduplicated code corpora across multiple programming languages, while CodeSearchNet (Husain et al., 2019) and CodeXGLUE (Lu et al., 2021) support retrieval, summarization, and translation tasks with instruction-style prompts. CodeForces-CoTs (Penedo et al., 2025) incorporates chain-of-thought annotations for programming tasks, supporting more explainable and step-wise code generation. These code-specific datasets play a vital role in developing FedLLMs that can support privacy-sensitive, domain-specific programming environments. A comparative summary of these code-related datasets is presented in Table 9.
@@ 5.1.5 Math Domain Instruction Fine-Tuning Datasets
For English-language datasets, GSM8K (Cobbe et al., 2021b) is a widely used benchmark for grade-school math word problems with detailed rationales. CoT-GSM8K [^16] https://huggingface.co/datasets/Dahoas/cot__\__gsm8k augments it with chain-of-thought explanations to support intermediate reasoning steps. Datasets such as MathInstruct (Yue et al., 2023c), MetaMathQA (Yu et al., 2023b), OpenR1-Math-220k (Allal et al., 2025), and Orca-Math Word Problems 200k (Mitra et al., 2024) provide high-quality, instruction-based math problems covering arithmetic, algebra, and word problem solving. The Hendrycks MATH Benchmark (Hendrycks et al., 2021c) and DeepMind Mathematics [^17] https://huggingface.co/datasets/di-zhang-fdu/DeepMind__\__Mathematics__\__QA Dataset offer more advanced, competition-style problems suitable for evaluating formal mathematical reasoning. Recent datasets like DAPO-Math-17k (Yu et al., 2025), Big-Math-RL-Verified (Albalak et al., 2025), and OpenMathInstruct-1 (Toshniwal et al., 2024) integrate reinforcement signals, verifiable proofs, or multi-step derivations, further pushing the boundaries of instruction-aligned mathematical LLMs.
@@ 5.1.6 Legal Domain Instruction Fine-Tuning Datasets
For English-language instruction tuning, Legal-QA-v1202020https://huggingface.co/datasets/dzunggg/legal-qa-v1 provides question–answer pairs covering legal concepts and procedures across various subfields of law. Pile of Law (Gao et al., 2020) is a large-scale corpus of U.S. legal documents—including court opinions, contracts, and regulations—that supports open-ended legal instruction tuning. CUAD (Hendrycks et al., 2021b) focuses on contract understanding, with annotated question–answer pairs tailored for clause extraction and risk analysis. LEDGAR [^21] https://huggingface.co/datasets/coastalchp/ledgar contains a large set of contractual clauses categorized into fine-grained legal functions, useful for classification and retrieval tasks in instruction-based formats.
For Chinese-language legal modeling, DISC-Law-SFT (Yue et al., 2023b) offers supervised instruction–response pairs across a wide spectrum of Chinese legal domains, including civil law, criminal law, and administrative law. Law-GPT-zh [^22] https://huggingface.co/datasets/sentence-transformers/law-gpt consolidates multiple sources of Chinese legal texts into an instruction tuning format to support legal consultation and statutory reasoning. Lawyer LLaMA Data (Huang et al., 2023a) further augments legal dialogue capabilities with simulated lawyer–client conversations, making it well-suited for on-device legal assistants in federated environments. A comparative overview of these law-specific instruction tuning datasets is provided in Table 9.
@@ 5.2 Evaluation Benchmarks
Table 10: Overview of general-domain evaluation benchmarks. Benchmarks specifically designed for long-context evaluation are highlighted. For each benchmark, we summarize its primary evaluation objective and the main evaluation criteria.
Benchmark Domain Evaluation Objective Main Evaluation Criteria
MMLU General Evaluate multitask language understanding across 57 subjects Multitask accuracy
BIG-bench General Evaluate advanced reasoning capabilities Model performance and calibration
DROP General Evaluate discrete reasoning over paragraphs Exact match and F1 score
CRASS General Evaluate counterfactual reasoning ability in LLMs Multiple-choice accuracy
ARC General Assess science reasoning at grade-school level Multiple-choice accuracy
AGIEval General Evaluate foundation models on human-centric standardized exam tasks Multi-task accuracy across disciplines
M3Exam General Evaluate multilingual, multimodal, and multilevel reasoning across real exam questions Multiple-choice accuracy
SCIBENCH General Evaluate college-level scientific problem-solving in math, physics, and chemistry Open-ended accuracy and skill-specific error attribution
Vicuna Evaluation General Evaluate instruction-following quality in chat settings Human/GPT-4 preference comparison accuracy
MT-Bench General Evaluate multi-turn conversational and instruction-following capabilities Winrate judged by GPT-4
AlpacaEval General Evaluate instruction-following performance via LLM-based auto-annotation Length-controlled win rate correlated with human preference accuracy
Chatbot Arena General Evaluate LLMs via human-voted battles using Elo rating Elo score and head-to-head win rate
PandaLM General Evaluate instruction-following quality and hyperparameter impact Winrate judged by PandaLM
HellaSwag General Evaluate commonsense inference by selecting the most plausible continuation Multiple-choice accuracy
TruthfulQA General Evaluate LLM truthfulness and avoidance of imitative falsehoods Truthfulness rate and MC accuracy
ScienceQA General Evaluate multimodal scientific reasoning and explanation generation Multiple-choice accuracy and explanation quality
Chain-of-Thought Hub General Evaluate LLMs’ multi-step reasoning across diverse domains using CoT prompting Few-shot CoT accuracy
NeuLR General Evaluate deductive, inductive, and abductive reasoning capabilities of LLMs Multi-dimensional evaluation (e.g., accuracy, self-awareness)
ALCUNA General Evaluate LLMs’ ability to comprehend and reason over novel knowledge Accuracy on 84,351 queries
LMExamQA General Evaluate LLMs’ performance across knowledge recall, understanding, and analysis Accuracy on 10,090 questions
SocKET General Evaluate LLMs’ sociability and understanding of social knowledge Accuracy and other metrics
Choice-75 General Evaluate LLMs’ decision reasoning ability within scripted scenarios Accuracy on binary multiple-choice questions
HELM General Evaluate LMs via multi-metric scenarios Composite performance ranking (normalized scores across metrics)
OpenLLM General Evaluate open-style reasoning across multiple benchmarks Normalized accuracy aggregate across tasks
BOSS General Evaluate out-of-distribution robustness across NLP tasks OOD accuracy drop and ID–OOD performance correlation
GLUE-X General Evaluate out-of-distribution robustness Average OOD accuracy drop relative to ID performance
PromptBench General Evaluate robustness, prompt-engineering and dynamic evaluation capabilities Adversarial success rate, robustness, prompt-variant performance
DynaBench General Evaluate robustness via dynamic, human-in-loop adversarial data collection Model error rate on human-crafted challenge examples
KoLA General Evaluate world knowledge across 19 evolving tasks self-contrast calibration
CELLO General Evaluate following complex, real-world instructions with multiple constraints Multi-criteria compliance rate judged by automated metrics
LLMEval General Evaluate LLM evaluators’ ability to assess instruction-following quality Meta-evaluator accuracy
Xiezhi General Evaluate holistic domain knowledge across 516 disciplines Mean Reciprocal Rank (MRR)
C-Eval General Evaluate Chinese domain knowledge and reasoning across 52 disciplines Multiple-choice accuracy
BELLE-eval General Evaluate Chinese instruction-following and multi-skill capabilities GPT-4 adjudicated win-rate and per-task quality scores
SuperCLUE General Evaluate Chinese instruction-following with human preference alignment GPT-4 adjudicated win-rate and Elo scores
M3KE General Evaluate Chinese LLMs’ knowledge across 71 disciplines and 4 education levels Multitask accuracy in zero- and few-shot settings
BayLing-80 General Evaluate cross-lingual and conversational capabilities GPT-4 adjudicated win-rate
MMCU General Evaluate multitask Chinese understanding across different domains Multitask accuracy (zero-/few-shot)
C-CLUE General Evaluate classical Chinese NER and relation extraction capabilities Weighted F1 score for NER and RE tasks
LongBench General (long-context) Evaluate bilingual long-context understanding across multiple real-world tasks Accuracy and generation quality
L-Eval General (long-context) Evaluate long-context reasoning and task performance over documents up to 60K tokens Multi-metric assessment
InfinityBench General (long-context) Evaluate long-context understanding across tasks up to 2M tokens in length Accuracy and task-specific metrics
Marathon General (long-context) Evaluate long-context understanding across multi-domain and multi-task settings Accuracy, F1, ROUGE-L, EM across tasks and context lengths
LongEval General (long-context) Evaluate LLMs’ effectiveness in long-context retrieval tasks Accuracy in broad-topic and fine-grained passage retrieval
BABILong General (long-context) Evaluate long-context reasoning in haystack-like settings Accuracy on long-context reasoning tasks
DetectiveQA General (long-context) Evaluate long-context reasoning through detective novel comprehension Instruction-following accuracy and winrate judged by PandaLM
NoCha General (long-context) Evaluate narrative comprehension and long-range coreference resolution in literary texts Exact match and accuracy
Loong General (long-context) Evaluate LLMs’ ability to perform multi-document reasoning and extended-context question answering Answer accuracy, context utilization rate, and document coverage score
TCELongBench General (long-context) Evaluate temporal reasoning over complex, long event narratives Temporal event ordering accuracy and event query answering accuracy
DENIAHL General (long-context) Evaluate the influence of in-context features on LLMs’ needle-in-a-haystack abilities Accuracy on long-context retrieval and reasoning tasks
LongMemEval General (long-context) Benchmark LLMs’ ability to retain and utilize long-term interactive memory across dialogue sessions Memory retention accuracy, retrieval success rate, and context utilization score
Long2RAG General (long-context) Evaluate long-form generation and retrieval grounding under extended input lengths Key point recall, LLM-based coherence, factuality, and answer relevance
L-CiteEval General (long-context) Evaluate LLMs’ ability to utilize citations and contextual evidence Citation recall, LLM-based factuality and faithfulness scoring
LIFBENCH General (long-context) Evaluate instruction-following performance and robustness in long-context scenarios Instruction-following accuracy and response stability
LongReason General (long-context) Evaluate long-context reasoning via synthetic context expansion Instruction-following accuracy and reasoning correctness
BAMBOO General (long-context) Evaluate long-text modeling capacity across diverse real-world tasks Instruction-following accuracy and winrate judged by PandaLM
ETHIC General (long-context) Evaluate instruction-following on long-context tasks with high information coverage Exact match, F1, and consistency/completeness
LooGLE General (long-context) Evaluate LLMs’ ability to understand and reason over long contexts Accuracy on inputs averaging 20K words
HELMET General (long-context) Evaluate instruction-following, retrieval, reasoning, summarization, and long-context understanding Task-specific automatic metrics and human evaluation
HoloBench General (long-context) Evaluate LLMs’ holistic reasoning ability over database-style textual inputs Execution accuracy and reasoning consistency on DB-style tasks
LOFT General (long-context) Evaluate whether long-context LLMs can replace retrieval-augmented methods in complex tasks EM/F1, SQL exec. accuracy, LLM-based factuality and reasoning coherence
Lv-Eval General (long-context) Evaluate long-context comprehension across five input length levels (up to 256k tokens) Exact match, ROUGE-L, LLM-based factuality and relevance scoring
ManyICLBench General (long-context) Evaluate many-shot in-context learning capabilities under extended context lengths Average accuracy across tasks under different context lengths
ZeroSCROLLS General (long-context) Evaluate zero-shot inference capabilities of LLMs on diverse long-text tasks Accuracy on inputs averaging 10K words
LongICLBench General (long-context) Evaluate long-context LLMs’ capability in in-context learning under extended input lengths Winrate judged by PandaLM
LIBRA General (long-context) Evaluate long-context understanding and instruction-following in Russian Accuracy, BLEU, LLM-based faithfulness and coherence scoring
@@ 5.2.1 General Evaluation Benchmarks
Several benchmarks target more specialized reasoning abilities: ScienceQA (Lu et al., 2022a) evaluates multimodal scientific question answering; Chain-of-Thought Hub (Fu et al., 2023) focuses on step-wise reasoning; NeuLR (Xu et al., 2023c) benchmarks deductive, inductive, and abductive reasoning; ALCUNA (Yin et al., 2023) assesses generalization to novel knowledge; LMExamQA (Bai et al., 2023c) tests models on recall, understanding, and analysis across over academic questions. Benchmarks like SocKET (Choi et al., 2023) and Choice-75 (Hou et al., 2023) address social knowledge and decision-making in scripted scenarios, respectively. Broad-scoped platforms such as HELM (Liang et al., 2022) and OpenLLM [^24] https://huggingface.co/spaces/open-llm-leaderboard/open__\__llm__\__leaderboard integrate multiple datasets and offer normalized aggregate scores across diverse metrics and tasks.
@@
(iv) Long-context understanding. Long-context reasoning is critical for FedLLM applications involving document-intensive tasks, extended dialogue, and memory retention. Benchmarks such as LongBench (Bai et al., 2023b), L-Eval (An et al., 2023), InfinityBench (Zhang et al., 2024h), Marathon (Zhang et al., 2023b), LongEval (Li et al., 2023a), and BABILong (Kuratov et al., 2024) form the backbone of this category, measuring comprehension, retrieval, and reasoning over contexts up to 2 million tokens. Further specialized benchmarks include: 1) Narrative and temporal reasoning: DetectiveQA (Xu et al., 2024d), NoCha (Karpinska et al., 2024), Loong (Wang et al., 2024b) and TCELongBench (Zhang et al., 2024k) focus on long-range narrative understanding and event-based temporal reasoning in complex textual sequences. 2) Retrieval and memory evaluation: DENIAHL (Dai et al., 2024), LongMemEval (Wu et al., 2024a), Long2RAG (Qi et al., 2024b), and L-CiteEval (Tang et al., 2024a) assess in-context feature sensitivity, long-term memory utilization, retrieval grounding, and the model’s ability to incorporate external citations.

### v3

@@ 5.1.3 Medical Domain Instruction Fine-Tuning Datasets
For English-language datasets, ChatDoctor-200K (Li et al., 2023i) and ChatDoctor-HealthCareMagic-100k (Li et al., 2023i) contain medical dialogues derived from professional consultation platforms, facilitating multi-turn reasoning and symptom analysis. The Medical Meadow (Han et al., 2023) suite offers curated datasets from sources such as CORD-19999https://huggingface.co/datasets/medalpaca/medical_\_meadow_\_cord19 and MedQA [^10] https://huggingface.co/datasets/medalpaca/medical_\_meadow_\_medqa, targeting tasks like biomedical question answering, literature summarization, and clinical fact verification. Additionally, HealthCareMagic-100k-en [^11] https://huggingface.co/datasets/wangrongsheng/HealthCareMagic-100k-en supports English patient–doctor interactions across various medical specialties.
For Chinese medical instruction tuning, a rich set of datasets has been developed to address the unique linguistic and clinical characteristics of Chinese healthcare scenarios. Notable examples include ChatMed-Consult-Dataset (Zhu et al., 2023b) and CMtMedQA (Yang et al., 2024a), which focus on consultation-style QA pairs. DISC-Med-SFT (Bao et al., 2023), HuatuoGPT-SFT (Li et al., 2023d), and Huatuo26M-Lite [^12] https://huggingface.co/datasets/FreedomIntelligence/Huatuo26M-Lite offer large-scale instruction–response pairs in clinical medicine, public health, and disease treatment. Traditional Chinese Medicine (TCM) is also covered through datasets like ShenNong-TCM (Zhu & Wang, 2023), enabling LLMs to support specific diagnostic and treatment tasks.
@@
Table 10: A summary of representative instruction fine-tuning datasets in code, math, and law domains.
Dataset Language Construction Domain Description
CodeAlpaca English Model Code GPT-generated code instructions (Alpaca-style)
Code Instructions 120k Alpaca English Human + Model Code Code generation instructions (Alpaca-style)
CodeContests English Human Code Competitive programming problems
CommitPackFT English Human Code Filtered GitHub commits with high-quality messages
ToolBench English Human + Model Code Instructions for multi-tool API usage
CodeParrot English Human Code Deduplicated and filtered Python files from GitHub
The Stack v2 Dedup English Human Code Large-scale deduplicated source code
CodeSearchNet English Human Code Code-language pairs for retrieval and semantic search
CodeForces-CoTs English Human + Model Code 10k CodeForces problems with CoT (DeepSeek R1)
CodeXGLUE Code Refinement English Human Code Buggy and fixed Java functions for code refinement
GSM8K English Human Math 8.5k grade school math word problems
CoT-GSM8k English Human + Model Math GSM8K extended with CoT reasoning steps
MathInstruct English Human + Model Math CoT and Program-of-Thought rationales
MetaMathQA English Human + Model Math Question augmentations from GSM8K and MATH
OpenR1-Math-220k English Human + Model Math 220k math problems with CoT (DeepSeek R1)
Competition MATH English Human Math 12.5K high school competition math problems
DeepMind Mathematics Dataset English Synthetic Math Algorithmically generated math problems
OpenMathInstruct-1 English Human + Model Math 1.8M math problems with code-interpreter solutions
Orca-Math Word Problems 200k English Synthetic Math 200K synthetic grade school math word problems
DAPO-Math-17k English Human + Model Math 17K diverse math problems for reasoning
Big-Math-RL-Verified English Human + Model Math 251K math problems with verifiable answers for RL
BELLE-math-zh Chinese Human + Model Math Chinese elementary school math problems
MathInstruct-Chinese Chinese Model Math Chinese version of MathInstruct
Legal-QA-v1 English Human Law 3.7K legal QA pairs from legal forums
Pile of Law English Human Law Dataset for legal-domain tasks
CUAD English Human Law 26K expert-annotated legal contract QA
LEDGAR English Human Law 1.45M expert-labeled legal contract clauses
DISC-Law-SFT Chinese Human + Model Law Comprehensive instructions for diverse legal tasks
Law-GPT-zh Chinese Human Law Legal sentence pairs for embedding models
Lawyer LLaMA Chinese Human + Model Law Dataset for Chinese legal tasks
@@ 5.1.4 Code Domain Instruction Fine-Tuning Datasets
Representative datasets include CodeAlpaca (Chaudhary, 2023) and Code Instructions 120k Alpaca [^13] https://huggingface.co/datasets/iamtarun/code_\_instructions_\_120k_\_alpaca, which extend the Alpaca instruction format to software engineering tasks such as debugging, function generation, and refactoring. CodeContests (Li et al., 2022c) focuses on competitive programming tasks and provides instruction–response pairs related to algorithmic problem solving. CommitPackFT (Muennighoff et al., 2023) offers commit-message generation tasks based on source code diffs, reflecting real-world software maintenance scenarios. ToolBench (Qin et al., 2023a) is designed to help models learn tool-augmented code generation through instruction-following examples.
Several large-scale pretraining and fine-tuning datasets are also widely used for code instruction alignment. CodeParrot [^14] https://huggingface.co/datasets/codeparrot/codeparrot-clean and The Stack v2 Dedup (Kocetkov et al., 2022) provide diverse and deduplicated code corpora across multiple programming languages, while CodeSearchNet (Husain et al., 2019) and CodeXGLUE (Lu et al., 2021) support retrieval, summarization, and translation tasks with instruction-style prompts. CodeForces-CoTs (Penedo et al., 2025) incorporates chain-of-thought annotations for programming tasks, supporting more explainable and step-wise code generation. These code-specific datasets play a vital role in developing FedLLMs that can support privacy-sensitive, domain-specific programming environments. A comparative summary of these code-related datasets is presented in Table 10.
@@ 5.1.5 Math Domain Instruction Fine-Tuning Datasets
For English-language datasets, GSM8K (Cobbe et al., 2021) is a widely used benchmark for grade-school math word problems with detailed rationales. CoT-GSM8K [^15] https://huggingface.co/datasets/Dahoas/cot_\_gsm8k augments it with chain-of-thought explanations to support intermediate reasoning steps. Datasets such as MathInstruct (Yue et al., 2023b), MetaMathQA (Yu et al., 2023b), OpenR1-Math-220k (Allal et al., 2025), and Orca-Math Word Problems 200k (Mitra et al., 2024) provide high-quality, instruction-based math problems covering arithmetic, algebra, and word problem solving. The Competition MATH Benchmark (Hendrycks et al., 2021c) and DeepMind Mathematics [^16] https://huggingface.co/datasets/di-zhang-fdu/DeepMind_\_Mathematics_\_QA Dataset offer more advanced, competition-style problems suitable for evaluating formal mathematical reasoning. Recent datasets like DAPO-Math-17k (Yu et al., 2025), Big-Math-RL-Verified (Albalak et al., 2025), and OpenMathInstruct-1 (Toshniwal et al., 2024) integrate reinforcement signals, verifiable proofs, or multi-step derivations, further pushing the boundaries of instruction-aligned mathematical LLMs.
@@ 5.1.6 Legal Domain Instruction Fine-Tuning Datasets
For English-language instruction tuning, Legal-QA-v1191919https://huggingface.co/datasets/dzunggg/legal-qa-v1 provides question–answer pairs covering legal concepts and procedures across various subfields of law. Pile of Law (Gao et al., 2020) is a large-scale corpus of U.S. legal documents—including court opinions, contracts, and regulations—that supports open-ended legal instruction tuning. CUAD (Hendrycks et al., 2021b) focuses on contract understanding, with annotated question–answer pairs tailored for clause extraction and risk analysis. LEDGAR [^20] https://huggingface.co/datasets/coastalchp/ledgar contains a large set of contractual clauses categorized into fine-grained legal functions, useful for classification and retrieval tasks in instruction-based formats.
For Chinese-language legal modeling, DISC-Law-SFT (Yue et al., 2023a) offers supervised instruction–response pairs across a wide spectrum of Chinese legal domains, including civil law, criminal law, and administrative law. Law-GPT-zh [^21] https://huggingface.co/datasets/sentence-transformers/law-gpt consolidates multiple sources of Chinese legal texts into an instruction tuning format to support legal consultation and statutory reasoning. Lawyer LLaMA (Huang et al., 2023a) further augments legal dialogue capabilities with simulated lawyer–client conversations, making it well-suited for on-device legal assistants in federated environments. A comparative overview of these law-specific instruction tuning datasets is provided in Table 10.
@@ 5.2.1 General Evaluation Benchmarks
Table 11: Overview of general evaluation benchmarks. Abbreviations: EM (Exact Match), OOD (Out-of-Distribution), NER (Named Entity Recognition), ICL (In-Context Learning), RE (Relation Extraction), MRR (Mean Reciprocal Rank), ASR (Adversarial Success Rate).
Benchmark Domain Evaluation Objective Main Evaluation Criteria
MMLU Gen. Multitask understanding (57 subjects) Accuracy
BIG-bench Gen. Advanced reasoning capabilities EM, ROUGE, and Accuracy
DROP Gen. Discrete reasoning over paragraphs EM & F1
CRASS Gen. Counterfactual reasoning Accuracy
ARC Gen. Grade-school science reasoning Accuracy
AGIEval Gen. Performance on human-centric exams Accuracy
M3Exam Gen. Multilingual, multimodal, multilevel reasoning Accuracy
SCIBENCH Gen. College-level scientific problem-solving Accuracy & Error Attribution
Vicuna Evaluation Gen. Instruction-following quality Human/GPT-4 Preference
MT‑Bench Gen. Multi-turn conversation & instruction-following GPT-4 Win Rate
AlpacaEval Gen. Instruction-following (LLM auto-annotation) Win Rate
Chatbot Arena Gen. Human-preference Elo Score
PandaLM Gen. Instruction-following & hyperparameter impact PandaLM Win Rate
HellaSwag Gen. Commonsense inference (continuation selection) Accuracy
TruthfulQA Gen. Truthfulness & falsehood avoidance Truthfulness Rate & Accuracy
ScienceQA Gen. Multimodal scientific reasoning & explanation Accuracy & Explanation Quality
CoT Hub Gen. Multi-step reasoning via CoT Accuracy
NeuLR Gen. Deductive, inductive, and abductive reasoning Accuracy
ALCUNA Gen. Novel knowledge comprehension & reasoning Accuracy
LMExamQA Gen. Knowledge recall, understanding, and analysis Accuracy
SocKET Gen. Social knowledge understanding Accuracy
Choice-75 Gen. Decision reasoning (scripted scenarios) Accuracy
HELM Gen. Holistic evaluation (multi-metric scenarios) Composite Score
OpenLLM Gen. Open-style reasoning (multi-benchmark) Normalized Accuracy
BOSS Gen. OOD robustness OOD Accuracy Drop
GLUE‑X Gen. OOD robustness OOD Accuracy Drop
PromptBench Gen. Robustness & prompt-engineering ASR & Accuracy
DynaBench Gen. Robustness (dynamic human-in-loop) Error Rate
KoLA Gen. Evolving world knowledge (19 tasks) Self-Contrast Calibration
CELLO Gen. Complex instruction-following Accuracy & BLEU
LLMEval Gen. Meta-evaluation of LLM evaluators Meta-evaluator Agreement
Xiezhi Gen. Holistic domain knowledge (516 disciplines) MRR
C‑Eval Gen. Chinese domain knowledge & reasoning Accuracy
BELLE-eval Gen. Chinese instruction-following & multi-skill GPT-4 Win Rate
SuperCLUE Gen. Chinese instruction-following (human-aligned) GPT-4 Win Rate & Accuracy
M3KE Gen. Chinese knowledge (71 disciplines, 4 levels) Accuracy
BayLing-80 Gen. Cross-lingual & conversational capabilities GPT-4 Win Rate
MMCU Gen. Multitask Chinese understanding Accuracy
C‑CLUE Gen. Classical Chinese NER & RE Accuracy, Recall, and F1
Table 12: Overview of general long-context (LC) evaluation benchmarks.
Benchmark Domain Evaluation Objective Main Evaluation Criteria
LongBench Gen. (LC) Bilingual long-context understanding Accuracy, F1, and ROUGE
L-Eval Gen. (LC) Long-context reasoning (up to 60k) Accuracy & Win Rate
InfiniteBench Gen. (LC) Long-context understanding (up to 2M) Accuracy & ROUGE
Marathon Gen. (LC) Long-context understanding (multi-domain) Accuracy
LongEval Gen. (LC) Long-context retrieval Accuracy
BABILong Gen. (LC) Long-context reasoning (haystack) Accuracy
DetectiveQA Gen. (LC) Long-context reasoning Accuracy & GPT-4 Judge Score
NoCha Gen. (LC) Narrative comprehension Accuracy
Loong Gen. (LC) Document reasoning & extended-context QA GPT-4 Judge Score
TCELongBench Gen. (LC) Temporal reasoning over long narratives Accuracy
DENIAHL Gen. (LC) Context feature influence ROUGE & EM
LongMemEval Gen. (LC) Long-term interactive memory EM
Long2RAG Gen. (LC) Long-text RAG capabilities BLEU, ROUGE, and EM
L-CiteEval Gen. (LC) Citation & evidence utilization ROUGE & Accuracy
LIFBENCH Gen. (LC) Long-context instruction-following Accuracy, ROUGE, and Pass@k
LongReason Gen. (LC) Long-context reasoning Accuracy & ROUGE-L
BAMBOO Gen. (LC) Long-text modeling (diverse tasks) Accuracy, F1, and ROUGE
ETHIC Gen. (LC) Long-context understanding F1, Win Rate, and EM
LooGLE Gen. (LC) Long-context understanding and reasoning GPT-4 Judge Score & Auto Metric
HELMET Gen. (LC) Multi-skill long-context capabilities Task Metrics & Human Evaluation
HoloBench Gen. (LC) Holistic reasoning (database-style text) Accuracy
LOFT Gen. (LC) Long-context reasoning Accuracy, EM, and Recall
Lv-Eval Gen. (LC) Long-context comprehension (up to 256k) Keyword Recall & F1
ManyICLBench Gen. (LC) Many-shot ICL capabilities Accuracy
ZeroSCROLLS Gen. (LC) Zero-shot inference capabilities ROUGE & F1
LongICLBench Gen. (LC) Long-context ICL Accuracy, ROUGE-L, and Pass@k
LIBRA Gen. (LC) Russian long-context understanding Accuracy, Recall, and F1
@@
Several benchmarks target more specialized reasoning abilities: ScienceQA (Lu et al., 2022a) evaluates multimodal scientific question answering; Chain-of-Thought Hub (Fu et al., 2023) focuses on step-wise reasoning; NeuLR (Xu et al., 2023c) benchmarks deductive, inductive, and abductive reasoning; ALCUNA (Yin et al., 2023) assesses generalization to novel knowledge; LMExamQA (Bai et al., 2023c) tests models on recall, understanding, and analysis across over academic questions. Benchmarks like SocKET (Choi et al., 2023) and Choice-75 (Hou et al., 2023) address social knowledge and decision-making in scripted scenarios, respectively. Broad-scoped platforms such as HELM (Liang et al., 2022) and OpenLLM [^23] https://huggingface.co/spaces/open-llm-leaderboard/open_\_llm_\_leaderboard integrate multiple datasets and offer normalized aggregate scores across diverse metrics and tasks.
@@
(iv) Long-context understanding. Long-context reasoning is critical for FedLLM applications involving document-intensive tasks, extended dialogue, and memory retention. Benchmarks such as LongBench (Bai et al., 2023b), L-Eval (An et al., 2023), InfiniteBench (Zhang et al., 2024h), Marathon (Zhang et al., 2023b), LongEval (Li et al., 2023a), and BABILong (Kuratov et al., 2024) form the backbone of this category, measuring comprehension, retrieval, and reasoning over contexts up to 2 million tokens. Further specialized benchmarks include: 1) Narrative and temporal reasoning: DetectiveQA (Xu et al., 2024e), NoCha (Karpinska et al., 2024), Loong (Wang et al., 2024b) and TCELongBench (Zhang et al., 2024k) focus on long-range narrative understanding and event-based temporal reasoning in complex textual sequences. 2) Retrieval and memory evaluation: DENIAHL (Dai et al., 2024), LongMemEval (Wu et al., 2024a), Long2RAG (Qi et al., 2024b), and L-CiteEval (Tang et al., 2024a) assess in-context feature sensitivity, long-term memory utilization, retrieval grounding, and the model’s ability to incorporate external citations.